# HapticLLaMA [Uncaptioned image]: A Multimodal Sensory Language Model for Haptic Captioning

### v1

@@ 3.3.1 Haptic Tokenizer
This tokenizer (Figure LABEL:fig:tokenizer:a) is motivated by the importance of spectral frequency information in characterizing vibrations Jones and Sarter (2008); Bensma¨ıa et al. (2005). Unlike the time domain, where signals are represented by amplitude values over time, the frequency domain represents a signal as a sum of sinusoids at varying frequencies. The frequency-based tokenizer begins by converting the time-domain signal into the frequency domain via the Fast Fourier Transform (FFT), and subsequently discretizes the resulting frequency components through binning into variable-width intervals. Our proposed tokenizer divides the frequency range into logarithmically spaced bins that correspond to just-noticeable differences in human frequency perception Choi and Kuchenbecker (2013); Israr et al. (2006). Similarly, the amplitude range is segmented into normalized levels. The tokenizer then assigns a unique token (e.g., FREQ_3_AMP_2) to each frequency-amplitude pair, encoding the signal’s spectral content into a form interpretable by LLMs.
@@ 3.3.2 Input Format
[…]
@@ 3.4.1 Stage 1: Supervised Fine-Tuning
where $W_{*}^{\text{lora}}$ denotes the LoRA-adapted weight matrix, and $W_{*}$ refers to the original pretrained weight matrix (e.g., the query $Q$ or value $V$ projection weights in a transformer). The subscripts $*\in\{Q,V\}$ indicate the Query and Value projection layers, respectively, and only the matrices $A_{*}$ and $B_{*}$ are trainable parameters.
@@
[…]
@@ 3.4.2 Stage 2: Fine-Tuning via RLHF
Random 5.775.775.77 2.592.592.59 10.6210.6210.62 11.0711.0711.07
LLaMA (signal-agnostic) 9.399.399.39 2.712.712.71 20.6720.6720.67 23.8423.8423.84
GPT-4.5 (signal-agnostic) 11.5311.5311.53 4.284.284.28 26.0926.0926.09 28.2028.2028.20
Frequency tokens + LLaMA 28.4628.4628.46 7.487.487.48 29.4929.4929.49 35.2435.2435.24
EnCodec tokens + LLaMA 30.0730.0730.07 8.618.618.61 29.5029.5029.50 35.2635.2635.26
[…]
@@ 5 Experiments
We set the following baselines: (1) Random: For each vibration in the test set, we randomly select one caption from the candidate set as the caption for that signal, (2) Signal-agnostic: Since LLMs can generate fluent descriptions grounded in their pretraining and guided through prompting, we examine the capability of LLaMA3.2-3B and GPT-4.5  [^4] https://platform.openai.com/docs/models/gpt-4.5-preview in generating captions without receiving any signal in the input (see Appendix A for details), (3) Without LoRA Finetuning: Fine-tuning is disabled by removing LoRA from the Frequency-based and EnCodec-based HapticLLaMA models, while keeping haptic token training. (4) Without RLHF: Model performance with and without DPO fine-tuning from human feedback is reported.
@@
Following prior work on captioning Drossos et al. (2020); Cui et al. (2018), we evaluate haptic captions using BLEU-1 Papineni et al. (2002), BLEU-4 Papineni et al. (2002), ROUGE-L Lin (2004), and METEOR Banerjee and Lavie (2005) metrics. Due to the inherent ambiguity of haptic perception, each vibration in HapticCap is associated with 8–10 ground-truth reference captions. We evaluate each prediction using the reference captions and report the mean and standard deviation of the metrics to reflect performance and variability. Implementation details are provided in Appendix C.
5.1 Results
@@ 5.3 Human Evaluation
Figure LABEL:fig:heatmap_bar presents the human evaluation results of descriptions generated by HapticLLaMA in Stage 1 and Stage 2, respectively. We use the human ratings from VibRate as the evaluation results after Stage 1. After completing both training stages, we randomly select 20 vibration signals from the HapticCap test set and generate corresponding haptic captions using the final HapticLLaMA model. Two evaluators are instructed to provide final ratings on a 1–7 scale, using the same human rating setup as in Stage 1.
The results show a noticeable improvement in the quality of generated captions after Stage 2, with over 61% of ratings above 3.5. Specifically, there is a visible shift in the distribution toward higher rating intervals, indicating that more captions are perceived as relevant and accurate after Stage 2. For instance, the proportion of higher ratings increases across all intervals after Stage 2, with about 10% increase in ratings above 3.5 and a particularly notable rise of about 3% and 4% in the range [5, 6) and [6, 7), respectively. In contrast, the share of lower ratings, such as [1, 2) and [2, 3.5), decreases significantly by around 7% and 3%, respectively. These results indicate that the preference alignment introduced in Stage 2 notably improves the model’s ability to generate descriptions that are semantically clearer and perceptually aligned with human judgments.
@@ 5.4 Category-Specific Performance
Figure LABEL:fig:category_performance compares HapticLLaMA variants across sensory, emotional, and associative categories using (a) BLEU-4 scores (with additional METEOR results in Appendix D) and (b) average human ratings. Both illustrate that the application of DPO consistently improves performance in all categories, yielding gains in BLEU-4 and average human rating, respectively, in line with the observed improvements in overall performance.
As measured by BLEU-4, emotional captions yield the highest performance among the three categories across both HapticLLaMA variants, with the EnCodec-based model achieving the leading score of 41.2 in BLEU-4. This outcome may be attributed to the fact that emotional captions tend to show higher agreement among multiple human annotators Hu et al. (2025), thereby facilitating more accurate and coherent generation.
According to human evaluations, sensory captions receive the highest ratings across all HapticLLaMA variants and stages. The discrepancy between automatic metrics and human judgments may stem from the limitations of automated methods (e.g., BLEU), which focus on surface-level textual similarity, whereas human raters are more attuned to the qualitative aspects of haptic experience. Similar discrepancy between automated metrics and human judgments are reported in image captioning literature Elliott and Keller (2014); Kilickaya et al. (2016).
@@ 5.5 Case Study
Figure 8 illustrates a comparison between reference ground-truth (R) and generated (G) captions for two haptic signals from the test set.
For the left signal, we observe a continuous haptic vibration featuring two prominent vibration peaks in the middle. Compared with the reference captions, the generated ones effectively capture continuity, weak intensity, and rhythm of the vibration in the sensory category. For the emotion dimension, the generated caption also successfully reflect emotional responses (e.g., “nervous” $\leftrightarrow$ “anxious”). Regarding the associative category, HapticLLaMA broadens the range of associations from “motorboat engine” to “damp shoes”, both capturing dull, heavy sensations and demonstrating a broader and more diverse associative space. The second signal displays features that are clearly distinct from those of the first. The vibration is an intermittent sequence of pulses with pauses at regular intervals. In the sensory caption, HapticLLaMA captures the pulsing characteristic (“short hit” $\leftrightarrow$ “mild thumps in between”) and overall energy (“strong” $\leftrightarrow$ “energetic”). For the emotional aspect, HapticLLaMA aligns with the overall sentiment and enhances it (e.g., adding “evoked excitement”). In terms of associative categories, “ringing phone” and “alarm clock” sensations are logically similar, with both involving sequences of strong pulses in mobile phones.
@@ 5.6 Haptic Tokenizer Analysis
Table 2 presents the summary statistics of haptic tokens for two tokenizers: Frequency and EnCodec.
EnCodec-based tokenizer has a significantly larger haptic token vocabulary (1,024) compared to the Frequency-based tokenizer (278). This suggests that EnCodec can represent haptic signals with greater granularity and expressive power, consistent with the improved performance observed in the EnCodec-based HapticLLaMA. The Frequency-based tokenizer produces shorter sequences when representing a vibration signal, with an average length of 47.5 tokens, indicating a more compact representation. In contrast, EnCodec outputs a fixed-length sequence of 1,379 tokens for all signals, which may introduce redundancy but ensures uniform input sizes for downstream models. In summary, the Frequency-based tokenizer offers higher compression efficiency, making it suitable for lightweight or real-time applications but slightly inferior performance. On the other hand, EnCodec provides more detailed and consistent representations, which may be beneficial for tasks requiring richer signal understanding.
@@ Conclusion
We propose HapticLLaMA, the first multimodal sensory language model trained with two distinct haptic tokenizers for the haptic captioning task. HapticLLaMA formalizes haptic signals into sequences of discrete tokens through a haptic tokenizer and integrates the haptic tokens with text into a prompt template. Built on the LLaMA architecture, HapticLLaMA is trained in two stages: (1) supervised fine-tuning on the LLaMA architecture with LoRA, and (2) fine-tuning through reinforcement learning from human feedback (RLHF). Our evaluation, combining automated generation metrics (e.g., BLEU and METEOR) and human assessments, shows that HapticLLaMA can effectively perceive haptic vibration signals and provide substantial improvements over existing LLMs on the haptic captioning task. Our work advances haptic-language understanding by enabling large language models to interpret physical sensory data.
@@ Appendix C Implementation Details
All experiments are conducted on NVIDIA RTX A100 and RTX H100 GPUs. Due to limited computational resources, we adopt LLaMA3.2-3B as the backbone of HapticLLaMA, and integrate Low-Rank Adaptation (LoRA) into all query and value projection layers of the Transformer architecture to enable parameter-efficient fine-tuning. In Stage 1, the batch size is set to 4, with an overall learning rate of 3e-4. The model is trained with a generative loss. In Stage 2, we adopt the default setting of DPO trainer to optimize HapticLLaMA. The model is optimized the policy to align with human preferences through a simple classification objective, thereby improving caption quality.
@@ Appendix D More Experimental Results
Figure 9 presents the METEOR scores of EnCodec HapticLLaMA across the three categories. The trends across the three categories are consistent with those observed in BLEU (see Figure LABEL:fig:category_performance(a)).

### v2

@@ 3.3.1 Haptic Tokenizer
This tokenizer (Figure LABEL:fig:tokenizer:a) is motivated by the importance of spectral frequency information in characterizing vibrations Jones and Sarter (2008); Bensmaïa et al. (2005). Unlike the time domain, where signals are represented by amplitude values over time, the frequency domain represents a signal as a sum of sinusoids at varying frequencies. The frequency-based tokenizer begins by converting the time-domain signal into the frequency domain via the Fast Fourier Transform (FFT), and subsequently discretizes the resulting frequency components through binning into variable-width intervals. Our proposed tokenizer divides the frequency range into logarithmically spaced bins that correspond to just-noticeable differences in human frequency perception Choi and Kuchenbecker (2013); Israr et al. (2006). Similarly, the amplitude range is segmented into normalized levels. The tokenizer then assigns a unique token (e.g., FREQ_3_AMP_2) to each frequency-amplitude pair, encoding the signal’s spectral content into a form interpretable by LLMs.
@@
(a) Two haptic tokenizers to encode vibration signals into input representations suitable for LLaMA.
@@ 3.3.2 Input Format
After tokenization, the haptic tokens are added to the vocabulary of LLaMA tokenizer and their embeddings are updated in the learning process. Given an input tuple $\langle s,c\rangle$, where $s$ represents the signal token sequence generated by the haptic tokenizer and $c$ denotes the category given from $\{\text{sensory},\text{emotional},\text{associative}\}$. For the input of HapticLLaMA, we concatenate haptic, category, textual prompt, and special tokens (e.g., <EOS>) as a multimodal prompt $I$. We use the prompt II=“haptic signal: <haptic tokens>, its <category> description is: <caption>.”. During training, we append the human-written caption as reference. The auxiliary text tokens (e.g., “its”, “description”, and “is”) can be interpreted as prompts that guide the model’s output.
@@ 3.4.1 Stage 1: Supervised Fine-Tuning
where $W_{*}^{\text{lora}}$ denotes the LoRA-adapted weight matrix, and $W_{*}$ refers to the original pretrained weight matrix (e.g., the query $Q$ or value $V$ projection weights in a transformer). The subscripts $\in\{Q,V\}$ indicate the Query and Value projection layers, respectively, and only the matrices $A_{*}$ and $B_{*}$ are trainable parameters.
@@
4: Formalize input based on $S$, $p$, and ccinto $p_{1},\cdots,s_{i},\cdots,p_{i},\cdots,\text{<EOS>}]$.
@@ 3.4.2 Stage 2: Fine-Tuning via RLHF
Random $5.77$ $2.59$ $10.62$ $11.07$
LLaMA (signal-agnostic) $9.39$ $2.71$ $20.67$ $23.84$
GPT-4.5 (signal-agnostic) $11.53$ $4.28$ $26.09$ $28.20$
Frequency tokens + LLaMA $28.46$ $7.48$ $29.49$ $35.24$
AST+GPT-2 43.51 16. 47 35.61 44.84
EnCodec tokens + LLaMA $30.07$ $8.61$ $29.50$ $35.26$
Frequency tokens + LLaMA + LoRA $40.12{\scriptstyle\pm 5.1}$ $23.16{\scriptstyle\pm 4.3}$ $42.95{\scriptstyle\pm 3.4}$ $50.16{\scriptstyle\pm 2.1}$
EnCodec tokens + LLaMA + LoRA $41.54{\scriptstyle\pm 3.2}$ $24.28{\scriptstyle\pm 2.1}$ $44.71{\scriptstyle\pm 2.4}$ $54.03{\scriptstyle\pm 2.3}$
Frequency HapticLLaMA[Uncaptioned image]: Frequency tokens + LLaMA + LoRA + RLHF $49.43{\scriptstyle\pm 1.4}$ $30.86{\scriptstyle\pm 1.5}$ $48.74{\scriptstyle\pm 2.1}$ $58.95{\scriptstyle\pm 2.1}$
EnCodec HapticLLaMA[Uncaptioned image]: EnCodec tokens + LLaMA + LoRA + RLHF $51.36{\scriptstyle\pm 1.3}$ $32.06{\scriptstyle\pm 1.2}$ $49.56{\scriptstyle\pm 1.6}$ $59.98{\scriptstyle\pm 1.8}$
@@ 5 Experiments
We set the following baselines: (1) Random: For each vibration in the test set, we randomly select one caption from the candidate set as the caption for that signal, (2) Signal-agnostic: Since LLMs can generate fluent descriptions grounded in their pretraining and guided through prompting, we examine the capability of LLaMA3.2-3B and GPT-4.5  [^4] https://platform.openai.com/docs/models/gpt-4.5-preview in generating captions without receiving any signal in the input (see Appendix A for details), (3) AST + GPT-2: AST serves as the HapticEncoder (analogous to the CLIP encoder in ClipCap) to encode raw haptic signal into haptic features, and GPT-2 is used as the caption decoder (i.e., non-LLM sequence model) for generating haptic descriptions. (4) Without LoRA Finetuning: Fine-tuning is disabled by removing LoRA from the Frequency-based and EnCodec-based HapticLLaMA models, while keeping haptic token training. (5) Without RLHF: Model performance with and without DPO fine-tuning from human feedback is reported.
@@
Following prior work on captioning Drossos et al. (2020); Cui et al. (2018), we evaluate haptic captions using BLEU-1 Papineni et al. (2002), BLEU-4 Papineni et al. (2002), ROUGE-L Lin (2004), and METEOR Banerjee and Lavie (2005) metrics. Due to the inherent ambiguity of haptic perception, each vibration in HapticCap is associated with 8–10 ground-truth reference captions. We evaluate each prediction using the reference captions and report the mean and standard deviation of the metrics to reflect performance and variability. Human evaluations are the primary benchmark, while automated metrics offer a complementary but imperfect proxy for scalable evaluation and ablation. Implementation details are provided in Appendix C.
@@ 5.1 Performance on Automated Metrics
(a) The distribution of human ratings after Stage 1 and 2 training for captions generated by EnCodec-based HapticLLaMA. Higher ratings are shown in blue and lower ratings are in gray.
@@ 5.2 Human Evaluation
Figure 8(a) presents the human evaluation results of descriptions generated by HapticLLaMA in Stage 1 and Stage 2, respectively. We use the human ratings from VibRate as the evaluation results after Stage 1. After completing both training stages, we randomly select 20 vibration signals from the HapticCap test set and generate corresponding haptic captions using the final HapticLLaMA model. Two evaluators are instructed to provide final ratings on a 1–7 scale, using the same human rating setup as in Stage 1. The results show a noticeable improvement in the quality of generated captions after Stage 2, with over 61% of ratings above 3.5. Specifically, there is a visible shift in the distribution toward higher rating intervals, indicating that more captions are perceived as relevant and accurate after Stage 2. For instance, the proportion of higher ratings increases across all intervals after Stage 2, with about 10% increase in ratings above 3.5 and a particularly notable rise of about 3% and 4% in the range [5, 6) and [6, 7), respectively. In contrast, the share of lower ratings, such as [1, 2) and [2, 3.5), decreases significantly by around 7% and 3%, respectively. These results indicate that the preference alignment introduced in Stage 2 notably improves the model’s ability to generate descriptions that are semantically clearer and perceptually aligned with human judgments.
(b) EnCodec-based HapticLLaMA’s performance across sensory, emotional, and associative categories.
@@ 5.3 Category-Specific Performance
Figure 8(b) compares HapticLLaMA variants across sensory, emotional, and associative categories using (a) BLEU-4 scores (with additional METEOR results in Appendix D) and (b) average human ratings. Both illustrate that the application of DPO consistently improves performance in all categories, yielding gains in BLEU-4 and average human rating, respectively, in line with the observed improvements in overall performance.
As measured by BLEU-4, emotional captions yield the highest performance among the three categories across both HapticLLaMA variants, with the EnCodec-based model achieving the leading score of 41.2 in BLEU-4. This outcome may be attributed to the fact that emotional captions tend to show higher agreement among multiple human annotators Hu et al. (2025), thereby facilitating more accurate and coherent generation. According to human evaluations, sensory captions receive the highest ratings across all HapticLLaMA variants and stages. The discrepancy between automatic metrics and human judgments may stem from the limitations of automated methods (e.g., BLEU), which focus on surface-level textual similarity, whereas human raters are more attuned to the qualitative aspects of haptic experience. Similar discrepancy between automated metrics and human judgments are reported in image captioning literature Elliott and Keller (2014); Kilickaya et al. (2016).
@@ 5.4 Case Study
Figure 8 illustrates a comparison between reference ground-truth (R) and generated (G) captions for two haptic signals from the test set. For the left signal, we observe a continuous haptic vibration featuring two prominent vibration peaks in the middle. Compared with the reference captions, the generated ones effectively capture continuity, weak intensity, and rhythm of the vibration in the sensory category. For the emotion dimension, the generated caption also successfully reflect emotional responses (e.g., “nervous” $\leftrightarrow$ “anxious”). Regarding the associative category, HapticLLaMA broadens the range of associations from “motorboat engine” to “damp shoes”, both capturing dull, heavy sensations and demonstrating a broader and more diverse associative space. The second signal displays features that are clearly distinct from those of the first. The vibration is an intermittent sequence of pulses with pauses at regular intervals. In the sensory caption, HapticLLaMA captures the pulsing characteristic (“short hit” $\leftrightarrow$ “mild thumps in between”) and overall energy (“strong” $\leftrightarrow$ “energetic”). For the emotional aspect, HapticLLaMA aligns with the overall sentiment and enhances it (e.g., adding “evoked excitement”). In terms of associative categories, “ringing phone” and “alarm clock” sensations are logically similar, with both involving sequences of strong pulses in mobile phones.
@@ 5.5 Haptic Tokenizer Analysis
Table 2 presents the summary statistics of haptic tokens for two tokenizers: Frequency and EnCodec. EnCodec-based tokenizer has a significantly larger haptic token vocabulary (1,024) compared to the Frequency-based tokenizer (278). This suggests that EnCodec can represent haptic signals with greater granularity and expressive power, consistent with the improved performance observed in the EnCodec-based HapticLLaMA. The Frequency-based tokenizer produces shorter sequences when representing a vibration signal, with an average length of 47.5 tokens, indicating a more compact representation. In contrast, EnCodec outputs a fixed-length sequence of 1,379 tokens for all signals, which may introduce redundancy but ensures uniform input sizes for downstream models. In summary, the Frequency-based tokenizer offers higher compression efficiency, making it suitable for lightweight or real-time applications but slightly inferior performance. On the other hand, EnCodec provides more detailed and consistent representations, which may be beneficial for tasks requiring richer signal understanding.
@@ Conclusion
We propose HapticLLaMA, the first multimodal sensory language model trained with two distinct haptic tokenizers for haptic captioning task. Built on the LLaMA architecture, HapticLLaMA is trained in two stages: (1) supervised fine-tuning on the LLaMA architecture with LoRA, and (2) fine-tuning through reinforcement learning from human feedback (RLHF). Our evaluation shows that HapticLLaMA can effectively perceive haptic vibration signals and provide substantial improvements over existing LLMs on the haptic captioning task. Our work advances haptic-language understanding by enabling LLMs to interpret physical sensory data.
@@ Appendix C Implementation Details
All experiments are conducted on NVIDIA RTX A100 and RTX H100 GPUs. Due to limited computational resources, we adopt LLaMA3.2-3B as the backbone of HapticLLaMA, and integrate Low-Rank Adaptation (LoRA) into all query and value projection layers of the Transformer architecture to enable parameter-efficient fine-tuning. In Stage 1, the batch size is set to 4, with an overall learning rate of 3e-4. The model is trained with a generative loss. For the frequency-based tokenizer, the bin edges are constructed using geometric spacing (via np.geomspace) based on the minimum and maximum values of each signal.
In Stage 2, we adopt the default setting of DPO trainer to optimize HapticLLaMA. The model is optimized the policy to align with human preferences through a simple classification objective, thereby improving caption quality. To construct preference pairs from the VibRate ratings, we proceed as follows: (1) Collect human-rated captions. Each caption generated by the Stage-1 model receives a human rating on a 1–7 scale. (2) Split captions into positive and negative groups. We use a 3.5 threshold (the midpoint of the scale) to separate captions into:high-rated (positive) group: ratings $\geq$ 3.5 and low-rated (negative) group: ratings < 3.5. (3) Form preference pairs. For DPO, every caption in the high-rated group is treated as the preferred response. Each preferred caption is then paired with all captions in the low-rated group, which serve as the rejected responses. This creates a set of (preferred, rejected) preference pairs. Regarding the pairing strategy, pairing only low-rated captions would train the model merely to choose the “less bad” option, rather than to prefer truly good captions. In contrast, pairing high-rated (positive) with low-rated (negative) captions explicitly encourages the model to favor higher-quality outputs, providing a much stronger and more meaningful supervisory signal.
This results in 2,585 preference pairs: 1,280 sensory, 849 emotional, and 456 associative. Each vibration has four rated captions per description type, allowing up to four preference pairs per category. This setup prevents data explosion during DPO training.
In training with DPO objective, HapticLLaMA learns to assign: (1) higher likelihood to the preferred caption, and (2) lower likelihood to the rejected caption conditioned on the same haptic input. Two models are involved: the policy model, which is trained to learn human preferences with VibRate dataset, and the reference model, which is the supervised fine-tuned model (e.g., the first-stage SFT model) used as the baseline prior to preference optimization. We set $\beta=0.1$ for DPO training.
Figure 9: Token-level analysis of three cases (from left to right: Case 1, Case 2, and Case 3), where “F” denotes the caption generated by frequency-based HapticLLaMA and “E” denotes the caption generated by EnCodec-based HapticLLaMA.
@@ Appendix D More Experimental Results
D.1 More Performance of HapticLLaMA
Figure 10 presents the METEOR scores of EnCodec HapticLLaMA across the three categories. The trends across the three categories are consistent with those observed in BLEU (see Figure 8(b)(a)).
@@ D.1 More Performance of HapticLLaMA
We also ran a separate experiment and constructed preference pairs with a different approach by taking any two captions associated with the same signal , labeling the higher-rated caption as the accepted choice and the lower-rated one as the rejected choice (max six pairs per signal and description category). This creates a perfectly balanced dataset for each category, butleads to a noticeable performance drop on each category, as shown in Table 6.
Category BLEU-4 METEOR
Sensory 17.81 ($\downarrow$1.54) 43.16 ($\downarrow$3.25)
Emotional 38.98 ($\downarrow$2.23) 58.62 ($\downarrow$4.16)
Associative 21.04 ($\downarrow$1.27) 49.46 ($\downarrow$2.47)
Table 6: Performance comparison in terms of BLEU-4 and METEOR across different categories.
D.2 Tokenizer Analysis
We create three example signals: (1) a regular rhythmic pulse, (2) a continuous signal with a single frequency, and (3) a continuous signal with various frequencies ranging from 50Hz-500Hz. We then generated sensory, emotional, and associative captions using both Frequency HapticLLaMA and EnCodec HapticLLaMA, as shown in Figure 9.
For the rhythmic pulse (case 1), EnCodec HapticLLaMA reliably produced captions that explicitly mentioned temporal rhythm (e.g., “rhythmic pattern”), whereas Frequency HapticLLaMA tended to describe only overall intensity without capturing rhythmic structure. For the single-frequency signal (case 2), both models produced similar descriptions, indicating that the frequency-based tokenizer is sufficient when no complex temporal dynamics are present. For the multi-frequency signal (case 3), EnCodec HapticLLaMA more often referred to changing or sweeping vibrations, suggesting that the EnCodec tokenizer better captures time-varying spectral patterns.